OpenStudy (anonymous):

consider the sequence An=12An-1with a1=2 the closed form of An is A.An=2+17^n-1 B.An=17n+19 C.An=2(17)^n-1 D.An=17n+2

hartnn (hartnn):

can you find A2 ?? plug in n=2 in \( A_n=12A_{n-1}\) and use A1 = 2, to get A2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

so what did u get for A2 ?

OpenStudy (anonymous):

24

hartnn (hartnn):

thats correct. now find the common difference 'd' = 2nd term -1st term = A2-A1 = ... ?

OpenStudy (anonymous):

a2-a1=wouldnt it be 23?

hartnn (hartnn):

sorry, i meant we have r = common ratio = A2/A1 = 24/2 =12 now we can directly use the formula \(\Large a_n = a_1 r^{n-1}\) here a1 = 2, r =12 just plug in!

hartnn (hartnn):

bdw, none of the choices work

hartnn (hartnn):

either question is incorrect, or the choices

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The fibonacci sequence is defined recurively by the following a1=1 a2=1 an=an-1-an-2 A.true b.false

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

next term is not the difference, its the sum., the correct one is : \(a_n=a_{n-1}+a_{n-2}\)

OpenStudy (anonymous):

so its false